### v2

# Pay Less Attention to Function Words for Free Robustness of Vision-Language Models

Abstract
To address the trade-off between robustness and performance for robust VLM, we observe that function words could incur vulnerability of VLMs against cross-modal adversarial attacks, and propose Function-word De-Attention (FDA) accordingly to mitigate the impact of function words. Similar to differential amplifiers, our FDA calculates the original and the function-word cross-attention within attention heads, and differentially subtracts the latter from the former for more aligned and robust VLMs. Comprehensive experiments include 2 SOTA baselines under 6 different attacks on 2 downstream tasks, 3 datasets, and 3 models. Overall, our FDA yields an average 18/13/53% ASR drop with only 0.2/0.3/0.6% performance drops on the 3 tested models on retrieval, and a 90% ASR drop with a 0.3% performance gain on visual grounding. We demonstrate the scalability, generalization, and zero-shot performance of FDA experimentally, as well as in-depth ablation studies and analysis. Code will be made publicly available.
1 Introduction
Building robust vision-language models (VLMs) has gathered profound academic focus because of the necessity of defending VLMs against various adversarial attacks. To this end, many works [17, 13] have been proposed to enhance model robustness, purify perturbations, or detect potential adversaries. Among them, adversarial training (AT) shows superior performance in enhancing the robustness of VLMs. However, AT-based methods incur significant performance drops compared to vanilla models and high computational costs.
To resolve the trade-off mentioned above, we propose to enhance VLM robustness by further refining vision-language alignment (VLA). Rather than perturbing images during fine-tuning, we break texts into finer grains: content words, i.e., nouns/verbs, and function words, i.e., am/is/are. Specifically, we hypothesize that function words could incur vulnerability of VLMs against cross-modal adversarial attacks because of their ubiquity and lack of specificity. To verify our hypothesis, we record the white-box similarity between function-/content-words and images during targeted (image) attacks [^1] We tested on the 1k testset of Flickr30k retrieval dataset, using PGD attack with $\epsilon={4}/{255}$., and find that 80.3% of images show higher similarity scores towards the function words than content words after attacks, while 0% of the images exhibit this pattern before attacks. We also provide a visual demonstration using Grad-CAM[18] from a successful untargeted attack to demonstrate the impact of function words. As shown in Fig.1, removing all function words greatly mitigates the distractions from adversarial perturbations. Lastly, to qualitatively validate the impact of function words in adversarial examples, we record the performance variation after removing nouns, adjectives, verbs, and function words on both clean and adversarial examples for Image-to-Text/Text-to-Image Retrieval (T2IR/I2TR), as presented in Table.1. Results show that function words are the only words that reduce ASR without causing a significant performance drop. These results confirm our hypothesis on function words, implying that proper removal of function words could potentially defend VLMs against attacks.
Consequently, inspired by the setting of differential transformers [22] and differential amplifiers, we propose Function-word De-Attention (FDA) as the first method to build robust VLMs by refining vision-language alignment. Specifically, our FDA works by deploying a parallel pipeline on multi-attention heads within fusion-encoders, calculating the cross-attention between function words and the input images, i.e., distractions. We further softmax along the dimensions of visual and textual tokens to highlight the most misleading textual/visual tokens. Finally, we subtract the above distractions from the original attention for the output. To validate the effectiveness of FDA, we conduct comprehensive experiments on two SOTA baselines, 3 models, 2 tasks, 3 datasets, and 6 attacks. Overall, our FDA yields an average 18/13/53% ASR drop with only 0.2/0.3/0.6% performance drops on the 3 tested models for retrieval, and a 90% ASR drop with better clean performance on grounding. Our FDA is also verified to enhance the zero-shot performance of VLMs.
Table 1: Effectiveness of removing different words when testing on clean and adversarial examples. Adversarial examples use AutoAttack, and $\Delta_{ASR}$ is presented using the average results for all epsilons ($\epsilon=2,4$). $\uparrow$ $\downarrow$ indicates increased/decreased $\Delta_{ASR}$ (higher values preferred). Removing function words can lower ASR without significantly harming clean performance.
Removed Clean (R@1) ($\uparrow$) Avg ASR Drop
Words T2IR I2TR $\Delta_{ASR}$ ($\uparrow$)
N/A 95.90 85.60 -
NOUN 58.90 32.83 $\uparrow$ 25.27
ADJ 91.60 78.36 $\downarrow$   0.42
VERB 93.80 79.36 $\downarrow$   0.38
FUNC 94.30 81.04 $\uparrow$   0.54
Figure 1: Grad-CAM of attention maps of VLM under white-box untargeted attacks through perturbed images. The texts are given at the bottom of the figure, with function words highlighted. Left: The VLM correctly recognizes the female student on the clean image given the token her. Mid: The VLM is distracted by the adversarial perturbation and partially looks at the male coach. Right: The distraction is mitigated by simply applying masks to remove all function words: the VLM successfully ‘looks back at’ the female student.
Overall, our contributions are summarized as follows:
  • We identify that function words are distractions for vision-language alignment and subsequently propose Function-word De-Attentioning (FDA) to pay less attention to function words for more aligned vision-language models with free robustness.
  • We conduct comprehensive experiments on two SOTA baselines, 3 models, 2 tasks, and 3 datasets, under 6 attacks, and validate the effectiveness of FDA in enhancing robustness while preserving performance.
  • We provide in-depth ablation studies to show qualitative analysis, the insensitivity of our FDA towards hyperparameters, generalization across backbones, and enhancement on zero-shot performance.
2 Related Work
Adversarial attacks on vision-language models. In light of the advancement in VLMs, adversarial attacks on VLMs have also emerged to fool VLMs into incorrect or misleading outputs. Recent studies on white-box attacks [1] have exhibited impressive results. Besides, black-box attacks[26, 11, 23, 5, 20] have also demonstrated significant effectiveness against pre-trained VLMs through transferable attacks.
Adversarial Defense on vision-language models For defenses, adversarial training (AT) [16, 25, 19] has exhibited significant effectiveness in defending models against various adversarial attacks against classification, retrieval, etc. Several AT-based methods [17, 13] have demonstrated impressive robustness boost on CLIP models. However, AT is notoriously well-known for downgrading performance significantly due to the inclusion of adversarial examples into training. Although [17] proposed FARE to use the visual embeddings of vanilla models as supervision to balance the trade-off between clean performance and robustness, the performance drops remain considerably noticeable. Besides, the high computational costs also hinder broader applications in practice.
3 Methodology
In this section, we first provide a brief preliminary for the original calculation pipeline of cross-attention and introduce our Function-words De-Attention (FDA).
3.1 Preliminary
For a given textual/visual encoder $\mathcal{T}/\mathcal{V}$, input images $I$ and texts $T$ are fed into repsective encoders with their attention masks $\mathcal{M}_{T}/\mathcal{M}_{I}$ for the embeddings $\mathbf{\mathcal{F}}_{T},\mathbf{\mathcal{F}}_{V}\in\mathbb{R}^{d_{k}}$:
$\mathcal{F}_{T}=\mathcal{T}(T,\mathcal{M}_{T}),\quad\mathcal{F}_{V}=\mathcal{V}(I,\mathcal{M}_{I})$ (1)
Then, cross-attention scores are calculated by inputting these hidden states into the fusion encoder:
$Att^{L,H}=softmax\Big(\frac{Q\big(\mathbf{\mathcal{F}}_{T}\big)K(\mathcal{F}_{V}\big)}{\sqrt{d_{k}}},dim=-1\Big)\cdot V(\mathcal{F}_{V}\big)$ (2)
where $Q/K/V$ is the query/key/value layers, and $L,H$ is the index of layers and attention heads.
3.2 Function-word De-Attention (FDA)
Figure 2: Left: An illustration of our Function-word De-Attention (FDA) method. On the existing process of attention calculation, which uses $\mathcal{F}_{V}$ and $\mathcal{F}_{T}$, we add a parallel pipeline to calculate the attentions between function words $\mathcal{F}_{T_{f}}$ and the images $\mathcal{F}_{V}$. Afterwards, the function-attention passes a control gate $\mathcal{G}$ before entering the FDA module (triangle) differentially to subtract distractions as presented in Eq.6. Right: We speculate that attacks can easily cross the boundary for misalignments for less aligned models (top), and by removing function-word distractions, models can learn a robust embedding (bottom), preventing misalignments.
Built upon our previous observation, we hypothesize that function words are potential distractions in vision-language alignment. To remove such distractions, we propose Function-word De-Attention (FDA): we add a parallel pipeline upon the existing cross-attention calculation to specifically acquire the cross-attention between function words and the input images, namely the distraction, and then subtract them from the original attention. An illustration of our FDA is given in Fig.2. We first parallelly extract the features of all function words (denoted as $T_{f}$) within the input texts by masking all other tokens, excluding [CLS]:
$\mathcal{F}_{T_{f}}=\mathcal{T}(T,\mathcal{M}_{T_{f}}),\forall T_{f}\in\mathcal{D}$ (3)
where $M_{t_{f}}$ is the function word mask, and $\mathcal{D}$ is the function word dictionary. Here, we use a dictionary shortlisted from the stopwords list in [9]. Subsequently, we adopt a parallel pipeline to calculate function words’ attention scores:
${S}_{T_{f}}^{L,H}=\frac{Q(\mathbf{\mathcal{F}}_{T_{f}})K(\mathbf{\mathcal{F}}_{V})^{T}}{\sqrt{d_{k}}}$ (4)
With the function words attention scores, we then conduct softmax along the dimensions of visual tokens and textual tokens, respectively. In this way, we highlight the visual tokens with false activation under token words or the most misleading tokens with the largest visual activation:
$\tilde{A}tt_{t}^{L,H}=softmax\big(S_{T_{f}}^{L,H},dim=-1\big)V,$ (5)
$\tilde{A}tt_{v}^{L,H}=softmax\big(S_{T_{f}}^{L,H},dim=-2\big)V$
Afterwards, we subtract both distractions from $Att$ individually and take the minimum value as the final attention scores, with $\mathcal{G}$ being a control gate for automatically adjusting the subtraction.
$\hat{A}tt^{L,H}=min\big(Att^{L,H}-\mathcal{G}(\tilde{A}tt_{t}^{L,H}),Att^{L,H}-\mathcal{G}(\tilde{A}tt_{v}^{L,H})\big)$ (6)
Finally, after complete calculation of FDA, denoted as $FDA(\cdot)$, we concatenate the attention scores from all attention heads for outputs:
$\mathbf{\hat{A}tt}^{L,H}=$ (7)
[…]
FDA can be flexibly implemented on any number of fusion layers/attention heads, as each may specialize differently [6]. A general intuition is to remove these distractions in the early layers instead of the later ones to avoid possible ‘absorption’ of distractions, but not so exquisitely upfront that it may undermine the contextual integrity of the original inputs. We provide ablation and analysis in Sec.4.3.
4 Experiments
Vision-Language Tasks&Datasets. To thoroughly evaluate the performance and robustness of FDA, we incorporate several downstream tasks, including Text-to-Image/Image-to-Text Retrieval (T2IR/I2TR) and Visual Grounding (VG). For datasets, we use the Flickr30k[14] and MSCOCO [10] dataset for retrieval, and RefCOCO+ [24] for VG.
Models. For T2IR/I2TR, we test our method on the ALBEF [7], TCL [21], and BLIP[8], using 14M/14M/124M pretrained images, respectively. All models use the ViT-B/32 [3] as visual encoders and BERT [2] as textual encoders. Specifically, TCL shares the same backbones as ALBEF but uses a different training strategy (triplet contrastive learning), while BLIP uses a larger pre-trained encoder with an extra decoder. CLIP [15] is not included due to the absence of fusion encoders.
Baselines. As for baselines, we adopt the two SOTA methods for robust CLIP, i.e., TeCoA [13] and FARE [17], on all the models as adversarial fine-tuning baselines. To account for the robustness and accuracy trade-off, we lower the perturbation strength of each method to ensure similar clean performance as our FDA, such that TeCoA and FARE serve as a reference to compare robustness. Specifically, we use $\epsilon=1$ for TeCoA and FARE. For Text-to-Image/Image-to-Text Retrieval, we adversarially fine-tune using TeCoA and FARE for 4/1/1 epochs for ALBEF/TCL/BLIP. For Visual Grounding, we adversarially train models with both TeCoA and FARE for only 1 epoch, as both methods incur significant performance drops on ALBEFs.
Attacks. To thoroughly evaluate the robustness of our models, we test all models with three adversarial attacks and use the average of all attacks for robustness evaluations. Specifically, we use Projected Gradient Descent [12] and AutoAttack [1], denoted as PGD and APGD. As for the adaptive attack, we apply function word masks to the input texts for APGD to evade our FDA, denoted Masked APGD (MAPGD). All attacks are fully white-box, i.e., the attackers are aware of and can access the extra FDA operations. For each attack, we follow the settings in [13] and [17] to attack images with perturbation bounded by $l_{\infty}=\frac{2}{255},\frac{4}{255}$. Specifically, for targeted attacks in T2IR/I2TR, we apply a circular shift targeted to ensure non-overlapping unmatched targets. For targeted attacks in VG, we follow the settings of [4] and do not apply a patched attack.
Metrics. We use the common metric, Attack Success Rate (ASR), to indicate the efficacy of all adversarial attacks. For an ASR on a given model and the baseline model, denoted as $ASR_{M}/ASR_{B}$, respectively, we calculate the relative ASR change in percentage using $\Delta_{ASR}=\frac{ASR_{B}-ASR_{M}}{ASR_{B}}\times 100\%$. Consequently, a positive/larger $\Delta_{ASR}$ indicates improved/stronger robustness, while a negative/lower $\Delta_{ASR}$ implies decreased/weaker robustness, with 0% (100%) meaning no robustness gain (completely defended). Details are given in the Sec.A of the Appendix.
Implementation Details. Since our FDA parallelly computes distracted attention for subtraction, finetuning models with FDA is identical to downstream finetuning without extra modifications or parameters. Following the settings in [7], we finetune the model by 10 epochs and use the last-epoch model for all tasks/models. For the layer index, we use $L^{0-1}$ and $H^{0-5}$ for all models/tasks/datasets, with corresponding ablation studies on the selection of the layer and attention head indices for all models and tasks in Sec.4.3.
Table 2: Attack success rate (ASR) of PGD/APGD/MAPGD (masked APGD) against for Text-to-Image/Image-to-Text Retrieval (T2IR/I2TR) on Flickr30k and $CO$. Results are presented in percentage (%). $\uparrow$ $\downarrow$ indicates increased/decreased $\Delta_{ASR}$ (higher values preferred). $\dagger$ indicates higher performance than clean models. (Full results are given in Sec.B of the Appendix.) Our FDA consistently shows the best performance and overall robustness on ALL models.
Dataset VLM $l_{\infty}$ Defense Text-to-Image Retrieval Image-to-Text Retrieval ASR drop
Clean (R@1) ASR ($\downarrow$) Clean (R@1) ASR ($\downarrow$)
PGD APGD MAPGD PGD APGD MAPGD $\Delta_{ASR}\uparrow$
Flickr ALBEF $\nicefrac{{2}}{{255}}$ No Defense 95.90 3.38 14.68 65.88 85.60 0.71 14.98 58.85 -
TeCoA 91.20 2.56 19.39 73.12 81.44 0.55 17.45 61.30 $\downarrow$   3.02
FARE 91.10 2.46 17.29 70.15 81.48 0.55 16.55 65.90 $\uparrow$   5.36
FDA 95.60 3.37 12.44 58.66 85.50 0.35 12.55 51.35 $\uparrow$ 22.26
$\nicefrac{{4}}{{255}}$
No Defense 95.90 8.72 16.09 80.92 85.60 7.20 15.89 77.14 -
TeCoA 91.20 9.13 19.34 85.48 81.44 4.60 18.45 79.60 $\downarrow$   2.15
FARE 91.10 9.27 18.60 86.25 81.48 5.20 18.35 80.60 $\downarrow$   2.48
FDA 95.60 7.90 13.70 75.80 85.50 4.90 13.70 71.00 $\uparrow$ 14.69
TCL
$\nicefrac{{2}}{{255}}$
No Defense 94.90 10.29 70.55 66.66 84.02 4.11 65.58 60.79 -
TeCoA 92.10 11.08 66.31 59.11 80.40 4.10 70.80 46.85 $\uparrow$   2.78
FARE 91.70 11.72 67.47 60.98 78.22 4.60 61.25 47.85 $\uparrow$   1.09
FDA 94.40   8.52 48.38 68.48 83.82 3.30 44.50 57.50 $\uparrow$ 14.29
$\nicefrac{{4}}{{255}}$
No Defense 94.90 37.66 81.11 81.63 84.02 29.72 78.36 73.10 -
TeCoA 92.10 44.29 80.62 80.08 80.40 35.40 76.60 67.95 $\downarrow$   4.16
FARE 91.70 46.21 81.03 79.64 78.22 38.05 76.95 67.35 $\downarrow$   6.42
FDA 94.40 30.36 58.64 86.24 83.82 24.25 56.50 77.80 $\uparrow$ 13.55
BLIP
$\nicefrac{{2}}{{255}}$ No Defense 97.20 25.10 63.26 50.19 87.30 11.83 60.08 44.35 -
TeCoA 90.30 19.28 59.38 48.67 78.04 8.85 47.80 37.15 $\uparrow$ 15.70
FARE 89.70 20.24 66.53 54.92 77.72 10.00 58.00 46.65 $\uparrow$   3.09
FDA 96.50   7.66 18.96 40.98 86.84 5.50 13.75 35.00 $\uparrow$ 51.60
$\nicefrac{{4}}{{255}}$
No Defense 97.20 61.18 86.39 71.27 87.30 67.00 86.08 71.60 -
TeCoA 90.30 62.39 88.85 75.69 78.04 62.35 87.35 72.30 $\downarrow$   1.09
FARE 89.70 66.29 92.35 80.49 77.72 67.30 90.50 82.50 $\downarrow$   7.04
FDA 96.50 15.86 28.37 60.64 86.84 14.45 16.30 55.50 $\uparrow$ 56.36
COCO ALBEF $\nicefrac{{2}}{{255}}$ No Defense 77.60 0.95 11.01 30.47 60.70 0.35 8.86 19.40 -
TeCoA 68.04 0.72 18.56 34.23 53.07 0.15 13.05 18.89 $\uparrow$   2.87
FARE 69.28 0.26 22.68 32.71 53.58 0.02 14.59 16.76 $\uparrow$   0.53
[…]
$\nicefrac{{4}}{{255}}$
No Defense 77.60 4.71 14.48 51.20 60.70 2.41 12.18 36.17 -
TeCoA 68.04 1.57 25.69 59.90 53.07 0.36 20.57 40.37 $\downarrow$   3.25
FARE 69.28 1.40 32.34 63.45 53.58 0.35 24.35 39.61 $\downarrow$ 16.08
[…]
Table 3: Attack success rate (ASR) of PGD/APGD/MAPGD (masked APGD) against for Visual Grounding (VG) on RefCOCO+. Results are presented in percentage (%). $\uparrow$ $\downarrow$ indicates increased/decreased $\Delta_{ASR}$ (higher values preferred). $\dagger$ indicates higher performance than clean models. (Full results are given in Sec.B of the Appendix.) Our FDA consistently shows the best performance and overall robustness on ALL models.
$l_{\infty}$ Defense Clean (Acc) ASR on Test A Split ($\downarrow$) ASR on Test B Split ($\downarrow$) Avg ASR drop
Val_d Test A Test B PGD APGD MAPGD PGD APGD MAPGD $\Delta_{ASR}\uparrow$
$\nicefrac{{2}}{{255}}$ No Defense 58.50 65.90 46.30 6.70 11.16 11.16 6.07 7.08 7.42 -
TeCoA 57.20 64.70 45.00 6.81 7.72 8.01 3.39 6.28 6.10 $\uparrow$ 21.21
FARE 56.40 64.20 44.70 6.13 10.42 9.96 4.54 6.72 6.21 $\uparrow$ 12.08
[…]
$\nicefrac{{4}}{{255}}$ No Defense 58.50 65.90 46.30 7.89 11.16 11.75 4.39 8.06 8.06 -
TeCoA 57.20 64.70 45.00 6.57 8.17 8.46 3.56 6.10 6.44 $\uparrow$ 21.63
FARE 56.40 64.20 44.70 6.74 9.66 10.27 4.03 7.06 6.55 $\uparrow$ 13.28
[…]
4.1 Main Results
In this section, we compare the robustness of FDA and other baselines on T2IR/I2TR and VG tasks.
T2IR/I2TR. Results of T2IR/I2TR on ALBEF, TCL, and BLIP for all methods are given in Table.2. Overall on Flickr, our FDA consistently exhibits the best robustness with the best clean performance on all models over all other baselines, yielding a 22.26/14.69%, 14.29/13.55%, and 51.60/56.36% average ASR drop over all 3 attacks on ALBEF/TCL/BLIP for $\epsilon=2/4$, with a negligible 0.30/0.10%, 0.50/0.22%, and 0.70/46% performance drops in R@1 for T2IR/I2TR on each model, respectively. On MSCOCO, similar patterns exist as our FDA boosts the ASR drop by 9/14% for $\epsilon=2/4$ with a 0.1% clean performance boost.
i. Attack-wise, on Flickr, our FDA exhibits the best defense against PGD and MAPGD in 22 out of 24 results, leading TeCoA/FARE by 60/65% on the BLIP model, demonstrating the effectiveness of FDA in enhancing robustness against various attacks. For the strongest adaptive attack, MAPGD, our FDA maintains its lead over TeCoA and FARE on ABLEF and BLIP, with an average lead by over 10%. Although our FDA shows more vulnerability against APGD on the TCL model, it retains the best comprehensive robustness of the other two baselines, yielding a 10-20% lead for $\epsilon=2/4$. It is noticeable that TeCoA/FARE becomes ineffective for all attacks with $\epsilon=4$, while our FDA retains its effectiveness facing stronger attacks. Similar trends also exist on MSCOCO.
ii. Performance-wise, all baseline methods suffered from a performance drop by an average4/3/9% on ALBEF/TCL/BLIP. Nevertheless, our FDA only causes minor or little drops of less than 1% for all models, yielding a lead of TeCoA and FARE by approximately 4%, 3%, and %/7 on average, demonstrating the feasibility of paying less attention to function words for free robustness.
iii. Scalability-wise, we find that the effectiveness of FDA benefits significantly as the model scales: on ALBEF/TCL, which uses 14M pre-trained images, FDA enhances robustness of each model by roughly 15%; while on BLIP, which uses 124M pre-trained images, FDA achieves an impressive 54% overall increase in $\Delta_{ASR}$. We attribute the enhancement to the capability of the backbone model, which enables the encoders to better capture visual clues.
Grounding. Similar patterns persist for VG as shown in Table.3. Our FDA achieves almost complete defense for all attacks, yielding an over 90% ASR drop while performing better on clean examples than the vanilla model. Specifically, FDA shows 93.16/91.50% ASR drop for $\epsilon=2/4$. While TeCoA and FARE show comparative clean performance, they achieve 21.21/21.63% and 12.08/13.28% ASR drop, respectively, with an over 1% drop on clean examples. These results confirm the efficacy of FDA in enhancing robustness for similar/better clean performance.
Table 4: Robustness evaluations on ALBEF using FDA as a plug-and-play tool with TeCoA and FARE against targeted and untargeted attacks for Text-to-Image/Image-to-Text Retrieval. Results are averaged over T2IR and I2TR. Full results are provided in Sec.C of the Appendix. FDA consistently boosts clean performance and/or robustness against all attacks.
VLM Defense Clean (R@1) Average ASR $\nicefrac{{2}}{{255}}$ ($\downarrow$) Average ASR $\nicefrac{{4}}{{255}}$ ($\downarrow$) Avg ASR drop
T2IR I2TR PGD APGD MAPGD PGD APGD MAPGD Robustness ($\uparrow$)
ALBEF No Defense 95.90 85.60 72.67 68.13 63.19 94.71 83.81 81.75 -
TeCoA 92.30 81.40 75.84 64.09 61.65 97.49 81.41 82.69 $\uparrow$   0.47
TeCoA + FDA 92.50 81.86 75.52 63.22 60.44 97.57 80.84 82.01 $\uparrow$   1.31
FARE 91.20 80.76 69.87 48.18 44.00 96.43 75.79 75.79 $\uparrow$ 13.09
FARE + FDA 91.40 80.80 70.70 47.95 44.54 96.42 74.84 73.57 $\uparrow$ 13.46
BLIP No Defense 97.20 87.30 78.17 77.08 67.65 99.80 94.01 89.82 -
TeCoA 81.50 68.00 48.01 41.23 38.16 95.37 75.41 72.39 $\uparrow$ 28.23
TeCoA + FDA 80.40 67.78 43.80 38.20 35.63 94.26 72.20 69.67 $\uparrow$ 32.18
FARE 89.70 77.72 47.51 53.62 51.45 90.37 78.71 77.01 $\uparrow$ 22.36
FARE + FDA 89.80 77.72 45.07 49.54 46.97 89.96 76.11 74.15 $\uparrow$ 25.91
Table 5: Robustness evaluations of FDA as a plug-and-play tool with TeCoA and FARE against targeted and untargeted attacks for Visual Grounding. Results are averaged over Test-A and Test-B. Full results are provided in Sec.C of the Appendix. FDA consistently boosts clean performance and robustness against all attacks.
Defense Clean (Acc) Average ASR $\nicefrac{{2}}{{255}}$ ($\downarrow$) Average ASR $\nicefrac{{4}}{{255}}$ ($\downarrow$) Avg ASR drop
Val_d Test A Test B PGD APGD MAPGD PGD APGD MAPGD $\Delta_{ASR}\uparrow$
No Defense 58.50 65.90 46.30 27.46 20.06 19.82 32.33 23.48 23.31 -
TeCoA 57.20 64.70 45.00 9.67 12.68 12.80 9.64 16.22 16.43 $\uparrow$ 39.68
TeCoA + FDA 57.00 64.90 45.30 10.37 12.85 12.01 9.84 15.22 15.67 $\uparrow$ 40.30
FARE 56.40 64.20 44.70 10.56 14.45 14.49 10.79 16.70 16.97 $\uparrow$ 34.69
FARE + FDA 56.10 63.70 44.70 11.25 13.02 13.05 10.46 15.20 15.50 $\uparrow$ 39.35
4.2 Untargeted Attacks
Apart from targeted attacks, we further evaluate the robustness against untargeted attacks. Thus, we retrained all models using TeCoA/FARE and their combination with our FDA to validate the effectiveness of FDA in defending against untargeted attacks.
For T2IR/I2TR, as presented in Table.4. Overall, we find FDA consistently boosts the robustness of TeCoA and FARE for all untargeted attacks on all models. Specifically, the scalability of FDA also applies after combining with TeCoA/FARE: both methods benefit more from FDA on the larger backbone of BLIP, yielding a 4/3% robustness boost. Furthermore, we notice that FDA also boosts the clean performance of both methods on ALBEF considerably, besides the improvement in robustness. For VG, we observe identical patterns: implementing FDA yields a solid robustness gain. For example, FARE experiences a significant robustness enhancement regarding untargeted attacks by 5%. In sum, our FDA compatibility works with both TeCoA and FARE to further boost their robustness against untargeted attacks.
4.3 Ablation Study
We now provide comprehensive studies on untargeted attacks, hyperparameters of FDA: encoder, dictionary, layer/head, and zero-shot performance. (See full results in Sec.D of the Appendix.)
4.3.1 FDA v.s. Masking & Other DA
We start by providing comparisons of our FDA and fine-tuning models by directly masking function words. We also include content words and nouns for thorough evaluation. As for variant of FDA, we further compare our FDA with Adjective DA (ADA) and Determiner DA (DDA). Specifically, we choose determiners (DET) and adjectives because DET indicates using a small subset (i.e., a/an/the) of function words, while ADJ adopts a completely different set of words. Results are presented in Table.6. Note: We only test on PGD and APGD since MAPGD is not applicable for nouns and content words.
Table 6: Comparison bewteen fine-tuning ALBEF by directly removing content words (CONT), nouns (NOUN), function words (FUNC), Determiner DA (DDA), and Adjective DA (ADA). $\Delta_{ASR}$ is presented using the average results for PGD and APGD. De-Attention shows significan advantages over directly masking and de-attentioning other words.
Maksd Clean (R@1) ($\uparrow$) Avg ASR Drop
Words T2IR I2TR $\Delta_{ASR}$ ($\uparrow$)
N/A 95.90 85.60 -
CONT 21.50 11.10 -
NOUN 68.60 44.62 -
FUNC 94.00 80.86 $\uparrow$  1.56
DDA 95.60 85.42 $\uparrow$  9.28
ADA 95.50 85.38 $\uparrow$ 15.10
FDA 95.60 85.50 $\uparrow$ 23.07
4.3.2 FDA v.s. Adaptive Wwords selection
[…]
In summary, we find that the gained robustness of VLMs increased as the proportion of function words increased. While we cannot design an adaptive mechanism that perfectly aligns with using the function words dictionary, we find that while the vulnerability of VLMs does not necessarily come from low-similarity (or low semantic) words, there is an evident correlation between the percentage of function words and the gained robustness.
Table 7: Comparison between FDA and adaptive selection, i.e., using image-text similarity to choose less informative tokens. SIM-$\delta$$2\delta$ indicates using $\mu-\delta$ and $\mu-2\delta$ as the de-attention threshold, with $\mu,\delta$ being the mean and std of the text-image similarity. SIM-N refers to choosing the lowest N tokens, with N being the number of function words in the text. % of words means the percentage of function words in the selected ones. Results confirm the correlation between the proportion of function words and the gained robustness.
Defense % of Words Clean R@1 ($\uparrow$) ASR Drop $\Delta_{ASR}$ $(\uparrow)$
in Dictionary T2IR I2TR $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ Avg
SIM-N 25.95 95.90 85.50 2.44 12.81 $\uparrow$  7.62
SIM-$2\delta$ 74.53 95.60 85.32 9.38 13.86 $\uparrow$  8.39
SIM-$\delta$ 79.49 95.30 85.38 12.85 11.54 $\uparrow$ 12.41
FDA 100.00 95.90 85.50 27.61 18.53 $\uparrow$ 23.07
Table 8: Ablation studies on the encoders and dictionary of FDA. We use T2IR/I2TR for evaluation.
Defense Clean (R@1) Average ASR $\nicefrac{{2}}{{255}}$ ($\downarrow$) Average ASR $\nicefrac{{4}}{{255}}$ ($\downarrow$) Avg ASR drop
T2IR I2TR Avg PGD APGD MAPGD PGD APGD MAPGD $\Delta_{ASR}\uparrow$
w/o FDA 95.90 85.60 90.75  2.04 14.83 62.37  7.96 15.99 79.03 -
$\mathcal{T}$ 95.10 85.28 90.19  2.10 21.86 8.15 10.66 24.70 17.30 $\downarrow$   2.54
$\mathcal{T}$ & $\mathcal{H}$ 93.80 85.00 89.40  2.01 17.61 15.82  9.06 21.91 20.99 $\uparrow$ 15.61
$\mathcal{H}$ 95.60 85.50 90.55 1.86 12.50 55.00  6.40 13.70 73.12 $\uparrow$ 18.48
Full Dict 95.10 84.46 89.78  2.03 13.54 56.60  6.46 14.47 74.65 $\uparrow$   4.22
Shortlisted Dict 95.40 85.40 90.40 1.71 13.60 56.78  6.92 14.15 75.07 $\uparrow$   6.45
Table 9: Ablation studies on the layer/head index $L/H$ of FDA on Text-to-Image/Image-to-Text Retreival on ALBEF, TCL and BLIP. Results are averaged over T2IR/I2TR. Shallower layers/heads (smaller $L/H$) consistently outperform over others on retrieval tasks.
VLM Defense Clean (R@1) Average ASR $\nicefrac{{2}}{{255}}$ ($\downarrow$) Average ASR $\nicefrac{{4}}{{255}}$ ($\downarrow$) Avg ASR drop
T2IR I2TR Avg PGD APGD MAPGD PGD APGD MAPGD $\Delta_{ASR}\uparrow$
ALBEF w/o FDA 95.90 85.60 90.75  2.04 14.83 62.37  7.96 15.99 79.03 -
${L^{all}},H^{all}$ 95.50 85.54 90.52  2.06 14.82 60.98  7.82 16.15 80.70 $\uparrow$   1.56
$L^{all},H^{6-11}$ 95.00 84.96 89.98  2.31 17.76 65.95  7.91 19.46 83.70 $\downarrow$   8.70
$L^{all},H^{0-5}$ 95.40 85.40 90.40 1.71 13.60 56.78  6.92 14.15 75.07 $\uparrow$   6.45
$L^{0}\quad,H^{0-5}$ 95.60 85.50 90.55 1.86 12.50 55.00  6.40 13.70 73.12 $\uparrow$ 18.48
$L^{0-1},H^{0-5}$ 95.40 85.32 90.36 1.81 12.30 54.87  6.17 13.45 72.71 $\uparrow$ 16.91
TCL w/o FDA 94.90 84.02 89.64 7.20 68.07 62.37 33.69 79.74 79.03 -
$L^{all},H^{0-5}$ 94.10 83.98 89.04 6.17 54.34 65.59 29.24 66.24 84.47 $\uparrow$   8.52
$L^{0}\quad,H^{0-5}$ 94.40 83.82 89.11 6.06 46.44 62.99 27.31 57.57 82.02 $\uparrow$ 13.92
$L^{0-1},H^{0-5}$ 94.20 83.96 89.08 6.42 48.64 64.82 28.22 61.30 83.44 $\uparrow$ 11.14
BLIP w/o FDA 97.20 87.30 92.25 18.46 61.67 47.27 64.09 86.23 71.44 -
$L^{all},H^{0-5}$ 96.50 86.94 91.72 16.60 22.74 43.06 61.09 31.38 65.19 $\uparrow$ 26.46
$L^{0}\quad,H^{0-5}$ 96.80 86.86 91.83 6.43 17.41 39.79 15.85 23.51 59.32 $\uparrow$ 52.51
$L^{0-1},H^{0-5}$ 96.70 86.84 91.77 6.58 16.36 37.99 15.15 22.34 58.07 $\uparrow$ 53.98
Table 10: Ablation studies on the layer/head index $L/H$ of FDA on Visual Grounding Retrieval on ALBEF, TCL, and BLIP. Results are averaged over Test A/B splits.
Defense Clean (Acc) Average ASR $\nicefrac{{2}}{{255}}$ ($\downarrow$) Average ASR $\nicefrac{{4}}{{255}}$ ($\downarrow$) Avg ASR drop
Val_d Test A Test B Avg PGD APGD MAPGD PGD APGD MAPGD $\Delta_{ASR}\uparrow$
w/o FDA 58.50 65.90 46.30 56.90 6.38 9.12 9.29   6.14 9.61 9.90 -
$L^{all},H^{0-5}$ 57.90 65.80 46.40 56.70 1.02 2.06 2.27 0.77 2.38 2.22 $\uparrow$ 81.83
$L^{0}\quad,H^{0-5}$ 58.00 65.90 46.40 56.77 1.72 3.09 2.93 1.85 2.60 2.76 $\uparrow$ 71.43
$L^{0-1},H^{0-5}$ 58.10 66.80 46.10 57.00 0.59 0.87 0.73 0.92 0.72 0.88 $\uparrow$ 92.33
Table 11: Zero-shot performance by applying FDA as a plug-and-play tool on T2IR/I2TR on ALBEF/BLIP and VG on ALBEF. T2IR/I2TR uses R@1/5/10, while VG uses accuracies.
Tasks Models Method Avg Performance
ALBEF w/o FDA 92.01 -
$L^{0}$ 92.02 $\uparrow$   0.01
Retrieval
$L^{0-1}$ 92.41 $\uparrow$   0.40
$L^{all}$ 92.17 $\uparrow$   0.16
BLIP
w/o FDA 92.24 -
$L^{0}$ 92.19 $\downarrow$   0.05
$L^{0-1}$ 92.22 $\downarrow$   0.02
$L^{all}$ 92.71 $\uparrow$   0.47
VG ALBEF w/o FDA 53.12 -
$L^{0}$ 52.72 $\downarrow$   0.40
$L^{0-1}$ 52.68 $\downarrow$   0.44
$L^{all}$ 53.34 $\uparrow$   0.22
First of all, masking content words and nouns yields the largest performance drop, making it unviable for robustness evaluation. This aligns with the intuition that these words carry extensive semantic information crucial for VLM tasks. Furthermore, masking function words leads to evident performance drop ($\sim$3%) and brings negligible robustness ($\sim$1%).As for DDA, as a subset of FDA, it shows almost identical clean performance, with a significant drop in robustness, indicating insufficient de-attentioning. ADA also shows subpar performance compared with FDA. Overall, we find that FDA leads the clean and adversarial performance among other variants.
4.3.3 Hyperparameters
The implementation of FDA, especially the macro-hypaerparameters influencing where to implement, would largely impact the subsequent performance of models. We first provide relative ablation studies to help understand the mechanics and design of our FDA.
Encoders&Dictionary. We start by comparing three implementations: FDA on text encoders, fusion encoders, and both, denoted as $\mathcal{T}$, $\mathcal{H}$, and $\mathcal{T}\&\mathcal{H}$. As presented in the top rows of Table.8, we find that $\mathcal{T}$ performs the worst among all, indicating that an early subtraction is insufficient for removing such subtraction. Although $\mathcal{T}\&\mathcal{H}$ provides a significant robustness boost, it costs an evident 2% performance drop on performance, implying that subtraction on both encoders is too strong and potentially causes contextual distortion. $\mathcal{H}$ performs the best as it helps models concentrate while preserving the contextual meaning.
For the dictionary, we use the off-the-shelf stopwords dictionary in [9], containing 208 words/symbols, denoted as Full Dict. Furthermore, we use a shortlisted dictionary, by only using the most commonly used function words, containing 93 crucial function words, denoted as Shortlisted Dict. Both dictionary settings are trained with FDA $L^{a}ll$ to maximize their impacts on training. As presented in the lower row of Table.8, there are no significant performance gaps between the two settings, with Full Dict performing slightly worse regarding both clean and adversarial examples. We attribute the minor degradation to the length of the stopwords dictionary, which could unnecessarily skim words and distort the context. We provide the shortlisted dictionary in Sec.E of the Appendix.
Attention Head &Layer. We then investigate the index of the layers $L$ and attention heads $H$ for retrieval and grounding, as presented in Table.9 and Table.10. Specifically, we train a series models using FDA but using different $L$ and $H$: for layers, we use all, 0-1, and 0 layers, denoted as $L^{all},L^{0-1},L^{0}$; for attention heads, we use all heads, 1st half (0-5) and the second half (6-11), denoted as $H^{all},H^{0-5},H^{6-11}$. For T2IR/I2TR, as shown in Table.9, we find that the shallow implementations of FDA, i.e., $L^{0}/L^{0-1},H^{0-5}$ consistently yield the best performance on robustness on all models. Specifically, $L^{0},H^{0-5}$ constantly achieves the best clean performance, leading other counterparts by 0.1-0.2%.
We further test the leading 3 settings on retrieval tasks, i.e., $L^{all}/L^{0}/L^{0-1},H^{0-5}$ on VG. As shown in 10, we find the shallow $L^{0-1},H^{0-5}$ settings still top w.r.t. both adversarial and clean examples, leading other settings by 10-20%/0.3-0.4%, respectively.
Overall, while FDA behaves slightly differently across settings/tasks, its effectiveness remains solid and insensitive to the head/layer parameters.
4.3.4 Zero-shot Performance
Finally, we adopt the three settings of FDA without finetuning to evaluate the zero-shot performance on different tasks (T2IR/I2TR/VG) on ALBEF and BLIP ($H^{0-5}$ is omitted and unchanged for all FDA). Results are presented in Table.11. We find that $L^{all}$ performs the best for all tasks and all models. This not only suggests that $L^{all}$ serves as the most generalizable setting for multiple VL tasks and models, but also implies the feasibility of FDA for performance boost on zero-shot tasks.
4.4 Analysis and Visualization
We notice that the results of APGD and MAPGD somewhat worsen after adversarial finetuning, e.g., TeCoA and FARE on ALBEF, FARE on BLIP in Table.2, etc. As previously illustrated in Fig.2, defending against targeted attacks requires a more aligned vision-language embedding. Consequently, we hypothesize that such abnormality potentially originates from the disruption in vision-language alignment brought by adversarial noise for enhanced robustness.
Figure 3: Left: T-SNE of the vision-language embedding of vanilla VLM, FDA, FARE, and TeCoA. Our FDA is the most aligned model. Right: Comparison of text-image similarity for vanilla VLM versus VLM + FDA. Our FDA yields better alignment with larger similarities and smaller variances.
To validate our speculation, we visualize the vision-language distribution of ALBEF together with TeCoA, FARE, and FDA, as shown in the left graph Fig.3. From the left graph, we find that both FARE and TeCoA (left column) yield a severely disrupted embedding, where images and texts sparsely scatter away from each other. On the other hand, our FDA (lower right) has the most aligned cross-modal embedding, as all images and texts remain tightly aligned with each other. To numerically compare the alignment of FDA and the vanilla model, we record the top 200 average white-box text-image similarity scores. As shown in the right figure of Fig.3, applying FDA generates higher average text-image similarity scores and lower variations.
4.5 Limitaiton
Besides subtraction, FDA could be potentially improved through a modular or algorithmic approach for more refined removal. Furthermore, we did not implement FDA to fine-tune a larger VLM or verify the effectiveness of FDA using LoRa due to the hardware limitation. Finally, our FDA is designed for backbones with a fusion encoder and thus not directly implementable for CLIP and other similar backbones. However, we believe implementation on CLIP-like models would be a valuable exploration for future work.
5 Conclusion
We propose Function-word De-Attention (FDA) calculates the original and the function-word cross-attention within attention heads, and differentially subtracts the latter from the former for more aligned and robust VLMs. Specifically, we tested the FDA on 2 downstream tasks, 3 datasets, and 3 models, and evaluated all methods under 6 attacks. By comparing with existing SOTA defenses, our FDA shows superiority of FDA in boosting robustness.
Appendix
Appendix A Details for attacks and evaluation metrics
We first introduce the attacks and evaluation metrics for each VL task, including the scenarios where a targeted attack is considered successful and the corresponding metrics.
Text-to-Image/Image-to-Text Retrieval. For T2IR, a successful targeted attack is only when the manipulated images emerge in the Top 1/5 position given the targeted text queries; for I2TR, a successful attack is only when the targeted texts emerge in the Top 1/5 position given the manipulated images as the query. Consequently, the ASR of T2IR/I2TR would be the hit rate at the top 1/5, i.e., the probability of appearance in the top 1/5 position, denoted as ASR@1/5. In the main paper, we use the average of ASR@1/5 as the overall ASR. Untargeted attacks follow the identical setting of existing works, i.e., lowering the R@1/5 of the victim models.
Visual Grounding. For visual grounding, we choose to obfuscate the model by fooling it into recognizing other objects as the target, or, if there is only one object in the image, locating the position of the object incorrectly (top-left corner). A successful attack is when the IOU of the targeted bounding box and the model bounding box is larger than 0.5, i.e., the model locates the object within the targeted bounding box. As for untargeted attacks, we follow existing settings to lower the accuracy of the victim model and calculate the drops as ASR.
Appendix B Full results for targeted attacks
In this section, we provide full results for all targeted attacks on all models and tasks. Specifically, for T2IR and I2TR, results on ALBEF is given in Table.12 and Table.13, results on TCL is given in Table.14 and Table.15, and results on BLIP is given in Table.16 and Table.17, respectively. Targeted attacks for visual grounding on ALBEF are given in Table.18.
Table 12: ASR of white-box targeted attacks against Text-to-Image Retrieval on Flickr30k and $CO$. The model is ALBEF. Changes over unattacked values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the attack success rate of the adversarial image showing up in the top-1/5 position of the targeted text queries.
Dataset $l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 95.90 0.30 (+0.20) 7.50 (+6.50) 14.60 (+14.50) 15.70 (+14.70) 50.10 (+50.00) 81.90 (+80.90)
Flickr
$\nicefrac{{2}}{{255}}$
TeCoA 91.20 0.20 (+0.20) 5.30 (+4.90) 18.70 (+18.70) 20.40 (+20.00) 59.90 (+59.90) 86.40 (+86.00)
FARE 91.10 0.10 (+0.10) 5.10 (+4.80) 17.20 (+17.20) 17.80 (+17.30) 58.10 (+57.90) 80.50 (+80.00)
FDA-$L^{0}$ 95.60 0.10 (+0.10) 7.30 (+6.60) 12.10 (+12.10) 13.40 (+12.70) 43.60 (+43.60) 73.90 (+73.20)
FDA-$L^{0-1}$ 95.40 0.20 (+0.20) 6.90 (+6.10) 12.00 (+12.00) 12.80 (+12.00) 43.30 (+43.30) 73.60 (+72.80)
FDA-$L^{all}$ 95.40 0.40 (+0.40) 5.90 (+5.20) 12.80 (+12.80) 14.20 (+13.50) 43.50 (+43.50) 77.30 (+76.60)
No Defense 95.90 4.30 (+4.20) 14.10 (+13.10) 16.50 (+16.40) 16.60 (+15.60) 75.00 (+74.90) 87.00 (+86.00)
$\nicefrac{{4}}{{255}}$
TeCoA 91.20 3.90 (+3.90) 14.70 (+14.30) 19.40 (+19.40) 19.60 (+19.20) 81.00 (+81.00) 90.00 (+89.60)
FARE 91.10 4.00 (+4.00) 14.80 (+14.50) 18.80 (+18.80) 18.80 (+18.30) 79.70 (+79.70) 85.90 (+85.40)
FDA-$L^{0}$ 95.60 2.90 (+2.90) 13.50 (+12.80) 13.90 (+13.90) 14.10 (+13.40) 69.00 (+69.00) 82.40 (+81.70)
FDA-$L^{0-1}$ 95.40 3.00 (+3.00) 12.40 (+11.60) 13.90 (+13.90) 14.00 (+13.20) 68.10 (+68.10) 81.30 (+80.50)
FDA-$L^{all}$ 95.40 3.00 (+3.00) 13.50 (+12.80) 14.60 (+14.60) 14.80 (+14.10) 68.90 (+68.90) 84.10 (+83.40)
COCO No Defense 77.60 0.22 (+0.18) 1.80 (+1.72) 10.18 (+10.14) 11.94 (+11.86) 20.14 (+20.14) 40.88 (+0.80)
$\nicefrac{{2}}{{255}}$
TeCoA 68.04 0.10 (+0.08) 0.66 (+0.56) 16.42 (+16.40) 17.40 (+17.34) 24.52 (+24.50) 44.02 (+43.92)
FARE 69.28 0.08 (+0.06) 0.55 (+0.45) 19.64 (+19.62) 25.82 (+25.72) 21.76 (+21.74) 43.74 (+43.64)
FDA-$L^{0-1}$ 77.70 0.26 (+0.22) 1.58 (+1.46) 9.00 (+  8.98) 10.40 (+10.30) 18.28 (+18.26) 37.00 (+36.90)
No Defense 77.60 2.10 (+2.06) 7.44 (+7.36) 14.26 (+14.22) 14.80 (+14.72) 43.74 (+43.70) 58.72 (+58.64)
$\nicefrac{{4}}{{255}}$
TeCoA 68.04 0.52 (+0.50) 2.74 (+2.64) 25.00 (+24.98) 26.46 (+26.36) 53.56 (+53.54) 66.28 (+66.18)
FARE 69.28 0.40 (+0.38) 2.52 (+2.42) 30.94 (+30.92) 33.82 (+33.72) 54.46 (+54.44) 72.48 (+72.38)
FDA-$L^{0-1}$ 77.70 1.74 (+1.70) 6.06 (+5.94) 11.76 (+11.74) 12.08 (+11.98) 37.92 (+37.90) 51.98 (+51.88)
Table 13: ASR of white-box targeted attacks against Image-to-Text Retrieval on Flickr30k and $CO$. The model is ALBEF. Changes over unattacked values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the attack success rate of the targeted text queries showing up in the top-1/5 position of the adversarial image.
Dataset $l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 85.60 0.30 (+0.30) 1.10 (+1.10) 14.40 (+14.40) 15.40 (+15.40) 53.50 (+53.50) 63.50 (+63.50)
Flickr
$\nicefrac{{2}}{{255}}$
TeCoA 81.44 0.10 (+0.08) 1.00 (+1.00) 16.42 (+16.40) 17.40 (+17.34) 56.90 (+56.90) 65.70 (+65.70)
FARE 81.48 0.10 (+0.10) 1.00 (+1.00) 19.64 (+19.62) 25.82 (+25.72) 56.30 (+56.30) 65.90 (+65.90)
FDA-$L^{0}$ 85.50 0.10 (+0.10) 0.60 (+0.60) 12.30 (+12.30) 12.80 (+12.80) 46.50 (+46.50) 56.20 (+56.20)
FDA-$L^{0-1}$ 85.32 0.10 (+0.10) 0.80 (+0.80) 12.10 (+12.10) 12.50 (+12.50) 46.90 (+46.90) 55.90 (+55.90)
FDA-$L^{all}$ 85.40 0.20 (+0.20) 1.00 (+1.00) 13.50 (+13.50) 13.70 (+13.70) 48.50(+48.50) 58.00 (+58.00)
No Defense 85.60 4.50 (+4.50) 9.80 (+9.80) 15.70 (+15.70) 15.90 (+15.90) 74.40 (+74.40) 79.00 (+79.00)
$\nicefrac{{4}}{{255}}$
TeCoA 81.44 2.80 (+2.80) 6.40 (+6.40) 18.30 (+18.30) 18.60 (+18.60) 78.10 (+78.10) 81.10 (+81.10)
FARE 81.48 3.40 (+3.40) 7.00 (+7.00) 18.30 (+18.30) 18.40 (+18.40) 77.00 (+77.00) 80.60 (+80.60)
FDA-$L^{0}$ 85.50 3.30 (+3.30) 6.50 (+6.50) 13.70 (+13.70) 13.70 (+13.70) 68.80 (+68.80) 72.40 (+72.40)
FDA-$L^{0-1}$ 85.32 3.10 (+3.10) 6.90 (+6.90) 13.40 (+13.40) 13.50 (+13.50) 69.30 (+69.30) 72.30 (+72.30)
FDA-$L^{all}$ 85.40 4.00 (+4.00) 7.80 (+7.80) 14.10 (+14.10) 14.20 (+14.20) 72.20 (+72.20) 75.20 (+75.20)
COCO No Defense 60.70 0.22 (+0.22) 0.50 (+0.48) 7.68 (+7.68) 10.04 (+10.02) 14.64 (+14.64) 24.16 (+21.14)
$\nicefrac{{2}}{{255}}$
TeCoA 53.07 0.02 (+0.02) 0.12 (+0.12) 10.82 (+10.82) 13.58 (+13.54) 14.32 (+14.32) 33.74 (+33.74)
FARE 53.58 0.00 (+0.00) 0.06 (+0.04) 11.80 (+11.80) 17.40 (+17.38) 21.62 (+12.62) 20.92 (+20.90)
FDA-$L^{0-1}$ 60.63 0.16 (+0.10) 0.42 (+0.40) 7.14 (+  7.14) 12.34 (+12.32) 16.72 (+16.72) 26.66 (+26.64)
No Defense 60.70 1.46 (+1.46) 3.38 (+3.36) 11.11 (+11.10) 13.26 (+13.24) 50.10 (+50.00) 81.90 (+80.90)
$\nicefrac{{4}}{{255}}$
TeCoA 53.07 0.16 (+0.16) 0.56 (+0.56) 18.58 (+18.58) 25.56 (+25.56) 38.42 (+38.42) 52.00 (+51.96)
FARE 53.58 0.22 (+0.22) 0.50 (+0.48) 21.84 (+21.84) 26.88 (+26.86) 31.94 (+31.94) 47.30 (+47.28)
FDA-$L^{0-1}$ 60.63 1.20 (+1.20) 2.92 (+2.90) 9.88 (+  9.88) 11.28 (+11.26) 27.74(+27.74) 37.94 (+37.92)
Table 14: ASR of white-box targeted attacks against Text-to-Image Retrieval on Flickr30k. The model is TCL. Changes over unattacked values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the attack success rate of the adversarial image showing up in the top-1/5 position of the targeted text queries.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 94.90 2.10 (+2.10) 18.80 (+18.40) 66.00 (+66.00) 75.20 (+74.80) 50.90 (+50.90) 82.50 (+82.10)
$\nicefrac{{2}}{{255}}$
TeCoA 92.10 3.10 (+3.10) 19.30 (+19.00) 61.00 (+61.00) 71.70 (+71.40) 44.20 (+44.20) 74.10 (+73.80)
FARE 91.70 3.20 (+3.20) 20.40 (+20.20) 63.10 (+63.10) 71.90 (+71.70) 46.80 (+46.80) 75.20 (+75.00)
FDA-$L^{0}$ 94.40 1.90 (+1.90) 15.40 (+15.10) 44.90 (+44.90) 52.00 (+51.70) 52.40 (+52.40) 84.60 (+84.30)
FDA-$L^{0-1}$ 94.20 2.40 (+2.40) 17.00 (+16.60) 46.30 (+46.30) 55.80 (+55.00) 54.40 (+54.40) 86.70 (+85.90)
FDA-$L^{all}$ 94.10 2.30 (+2.30) 16.80 (+16.40) 50.80 (+50.80) 60.90 (+60.50) 54.90 (+54.90) 87.10 (+86.70)
No Defense 94.90 21.50 (+21.50) 54.00 (+53.60) 80.30 (+80.30) 82.00 (+81.60) 75.20 (+75.20) 88.10 (+87.70)
$\nicefrac{{4}}{{255}}$
TeCoA 92.10 27.80 (+27.80) 60.90 (+60.60) 79.70 (+79.70) 81.60 (+81.30) 74.10 (+74.10) 86.10 (+85.80)
FARE 91.70 30.20 (+30.20) 62.30 (+62.10) 80.30 (+80.30) 81.80 (+81.60) 73.20 (+73.20) 86.10 (+85.90)
FDA-$L^{0}$ 94.40 17.90 (+17.90) 43.00 (+42.70) 56.80 (+56.80) 60.60 (+60.30) 80.20 (+80.20) 92.30 (+92.00)
FDA-$L^{0-1}$ 94.20 18.80 (+18.80) 44.90 (+44.50) 60.20 (+60.20) 64.50 (+63.70) 80.40 (+80.40) 92.90 (+92.10)
FDA-$L^{all}$ 94.10 19.00 (+19.00) 44.90 (+44.50) 66.10 (+66.10) 69.10 (+68.70) 79.80 (+79.80) 94.50 (+94.10)
Table 15: ASR of white-box targeted attacks against Image-to-Text Retrieval on Flickr30k. The model is TCL. Changes over unattacked values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the attack success rate of the targeted text queries showing up in the top-1/5 position of the adversarial image.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 84.02 2.50 (+2.50) 5.70 (+5.70) 64.10 (+64.10) 66.80 (+66.80) 50.70 (+50.70) 58.90 (+58.90)
$\nicefrac{{2}}{{255}}$
TeCoA 80.40 2.10 (+2.10) 6.10 (+6.10) 69.60 (+69.60) 72.00 (+72.00) 65.50 (+65.50) 70.40 (+70.40)
FARE 78.22 3.10 (+3.10) 6.10 (+6.10) 59.70 (+59.70) 62.80 (+62.80) 65.20 (+65.20) 69.50 (+69.50)
FDA-$L^{0}$ 83.82 1.90 (+1.90) 5.30 (+5.30) 43.40 (+43.40) 45.60 (+45.60) 52.90 (+52.90) 62.10 (+62.10)
FDA-$L^{0-1}$ 83.96 1.80 (+1.80) 4.80 (+4.80) 45.30 (+45.30) 47.50 (+47.50) 53.90 (+53.90) 64.40 (+64.40)
FDA-$L^{all}$ 83.98 1.30 (+1.30) 4.60 (+4.60) 51.50 (+51.50) 54.30 (+54.30) 56.00 (+56.00) 64.40 (+64.40)
No Defense 84.02 24.60 (+24.60) 34.70 (+34.70) 78.80 (+77.80) 78.60 (+78.60) 70.40 (+70.40) 75.50 (+75.50)
$\nicefrac{{4}}{{255}}$
TeCoA 80.40 29.60 (+29.60) 34.70 (+34.70) 76.00 (+76.00) 77.20 (+77.20) 65.50 (+65.50) 70.40 (+70.40)
FARE 78.22 33.30 (+33.30) 39.50 (+39.50) 76.20 (+76.20) 77.70 (+77.70) 65.20 (+65.20) 69.50 (+69.50)
FDA-$L^{0}$ 83.82 20.00 (+20.00) 28.50 (+28.50) 56.10 (+56.10) 56.90 (+56.90) 75.50 (+75.50) 80.10 (+80.10)
FDA-$L^{0-1}$ 83.96 20.10 (+20.10) 29.30 (+29.30) 60.00 (+60.00) 60.80 (+60.80) 77.80 (+77.80) 82.70 (+82.70)
FDA-$L^{all}$ 83.98 22.20 (+22.20) 31.10 (+31.10) 64.40 (+64.40) 65.50 (+65.50) 79.50 (+79.50) 84.10 (+84.10)
Table 16: ASR of white-box targeted attacks against Text-to-Image Retrieval on Flickr30k. The model is BLIP. Changes over unattacked values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the attack success rate of the adversarial image showing up in the top-1/5 position of the targeted text queries.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 97.20 2.50 (+2.50) 46.10 (+46.10) 80.50 (+81.10) 75.20 (+74.80) 57.90 (+57.90) 84.70 (+84.30)
$\nicefrac{{2}}{{255}}$
TeCoA 81.50 4.30 (+4.30) 16.20 (+16.20) 19.30 (+19.30) 45.70 (+45.60) 14.80 (+14.80) 39.20 (+39.10)
FARE 79.40 1.00 (+1.00) 10.30 (+10.10) 12.90 (+12.90) 46.20 (+46.20)   9.70 (+  9.70) 38.00 (+37.90)
FDA-$L^{0}$ 96.80 3.10 (+3.10) 12.00 (+11.90) 13.60 (+13.60) 26.30 (+26.20) 24.20 (+24.20) 61.70 (+61.60)
FDA-$L^{0-1}$ 96.50 3.00 (+3.00) 12.40 (+12.30) 12.30 (+12.30) 25.60 (+25.70) 22.10 (+22.10) 59.90 (+59.80)
FDA-$L^{all}$ 96.50 3.20 (+3.20) 42.00 (+41.80) 16.00 (+16.00) 39.60 (+39.40) 24.50 (+24.50) 66.30 (+66.30)
No Defense 97.20 31.80 (+31.80) 90.60 (+90.20) 79.80 (+79.80) 93.00 (+92.60) 57.90 (+57.90) 84.70 (+84.30)
$\nicefrac{{4}}{{255}}$
TeCoA 81.50 46.20 (+46.20) 73.00 (+72.90) 60.40 (+60.40) 83.10 (+83.00) 49.50 (+49.50) 74.80 (+74.80)
FARE 79.40 23.90 (+23.90) 61.50 (+61.30) 42.30 (+42.30) 82.70 (+82.60) 33.90 (+33.90) 74.70 (+74.70)
FDA-$L^{0}$ 96.80 13.60 (+13.60) 19.80 (+19.70) 16.40 (+16.40) 43.60 (+43.50) 44.70 (+44.70) 78.50 (+78.40)
FDA-$L^{0-1}$ 96.50 13.20 (+13.20) 18.60 (+18.50) 15.10 (+15.10) 41.60 (+41.70) 43.40 (+43.40) 77.90 (+77.80)
FDA-$L^{all}$ 96.50 31.60 (+31.60) 86.80 (+86.60) 21.40 (+21.40) 62.00 (+61.80) 50.00 (+50.00) 81.90 (+81.70)
Table 17: ASR of white-box targeted attacks against Image-to-Text Retrieval on Flickr30k. The model is BLIP. Changes over unattacked values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the attack success rate of the targeted text queries showing up in the top-1/5 position of the adversarial image.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 87.30 7.00 (+7.00) 16.60 (+16.60) 54.90 (+54.90) 65.00 (+65.00) 37.60 (+37.60) 50.90 (+50.90)
$\nicefrac{{2}}{{255}}$
TeCoA 68.00 2.40 (+2.40)   5.90 (+  5.90) 13.90 (+13.90) 25.10 (+25.10)   9.30 (+  9.30) 19.80 (+19.80)
TeCoA + FDA-$L^{0-1}$ 67.78 2.20 (+2.20)   6.20 (+  6.20) 12.90 (+12.90) 25.40 (+25.40)   8.90 (+  8.90) 19.60 (+19.60)
FARE 65.64 0.40 (+0.40)   1.90 (+  1.90) 11.30 (+11.30) 19.60 (+19.60)   8.00 (+  8.00) 15.10 (+15.10)
FARE + FDA-$L^{0-1}$ 66.22 0.40 (+0.40)   2.10 (+  2.10)   9.40 (+  9.40) 17.50 (+17.50)   6.00 (+  6.00) 13.60 (+13.60)
FDA-$L^{0}$ 86.86 4.40 (+4.40)   5.30 (+  5.30) 14.10 (+14.10) 15.70 (+15.70) 31.70 (+31.70) 41.60 (+41.60)
FDA-$L^{0-1}$ 86.86 4.60 (+4.60)   4.80 (+  4.80) 13.10 (+13.10) 14.40 (+14.40) 30.60 (+30.60) 39.40 (+39.40)
FDA-$L^{all}$ 86.94 6.70 (+3.20) 14.60 (+14.60) 16.90 (+16.90) 18.60 (+18.60) 35.00 (+35.00) 46.50 (+46.50)
No Defense 84.02 58.80 (+58.80) 74.90 (+74.90) 83.70 (+83.70) 88.10 (+88.10) 67.00 (+67.00) 75.90 (+75.90)
$\nicefrac{{4}}{{255}}$
TeCoA 68.00 46.60 (+46.60) 58.10 (+58.10) 58.90 (+58.90) 69.00 (+69.00) 47.10 (+47.10) 60.10 (+60.10)
TeCoA + FDA-$L^{0-1}$ 67.78 44.50 (+44.50) 59.10 (+59.10) 59.10 (+59.10) 70.30 (+70.30) 47.40 (+47.40) 61.10 (+61.10)
FARE 65.64 23.90 (+23.90) 37.30 (+37.30) 45.70 (+45.70) 60.90 (+60.90) 34.00 (+34.00) 50.90 (+50.90)
FARE + FDA-$L^{0-1}$ 66.22 24.70 (+24.70) 37.10 (+37.10) 43.90 (+43.90) 58.60 (+58.60) 31.90 (+31.90) 48.20 (+48.20)
FDA-$L^{0}$ 86.86 14.80 (+14.80) 15.30 (+15.30) 16.80 (+16.80) 17.30 (+17.30) 53.40 (+53.40) 60.70 (+60.70)
FDA-$L^{0-1}$ 86.86 14.20 (+14.20) 14.70 (+14.70) 16.20 (+16.20) 16.40 (+16.40) 52.00 (+52.00) 59.00 (+59.00)
FDA-$L^{all}$ 86.94 55.20 (+52.20) 70.80 (+70.80) 21.80 (+21.80) 22.10 (+22.10) 60.20 (+60.20) 68.70 (+68.70)
Table 18: Attack success rate (ASR) of targeted PGD/APGD/MAPGD (masked APGD) against for Visual Grounding (VG) on RefCOCO+. All results are presented in percentage (%). Changes over unattacked values are presented in parentheses.
$l_{\infty}$ Defense Clean Performance Test A Split ($\downarrow$) Test B Split ($\downarrow$)
Val_d Test A Test B PGD APGD MAPGD PGD APGD MAPGD
$\nicefrac{{2}}{{255}}$ No Defense 58.50 65.90 46.30 16.40 (+6.00) 20.40 (+10.00) 20.40 (+10.00) 25.73 (+4.80) 26.53 (+5.60) 27.30 (+6.37)
TeCoA 57.20 64.70 45.00 17.87 (+6.00) 18.67 (+  6.80) 18.93 (+  7.06) 24.00 (+2.67) 26.27 (+4.94) 26.13 (+4.80)
FARE 56.40 64.20 44.70 18.40 (+5.33) 22.13 (+  9.06) 22.00 (+  8.93) 24.27 (+3.60) 26.00 (+5.33) 25.87 (+5.20)
FDA-$L^{0}$ 58.00 65.90 46.40 12.53 (+1.73) 14.40 (+  3.60) 14.40 (+  3.60) 20.80 (+1.20) 21.33 (+1.73) 21.07 (+1.47)
FDA-$L^{0-1}$ 58.10 66.80 46.10 12.67 (+1.20) 13.60 (+  2.13) 13.06 (+  1.59) 20.26 (-0.14) 19.87 (-0.53) 20.13 (-0.27)
FDA-$L^{all}$ 57.90 65.80 46.40 12.27 (+2.14) 12.93 (+  2.80) 13.30 (+  3.17) 20.53 (-0.27) 21.60 (+0.80) 21.60 (+0.80)
$\nicefrac{{4}}{{255}}$ No Defense 58.50 65.90 46.30 17.47 (+7.07) 20.40 (+10.00) 20.40 (+10.00) 24.40 (+3.47) 26.53 (+5.60) 27.30 (+6.37)
TeCoA 57.20 64.70 45.00 18.00 (+6.13) 19.07 (+  7.20) 19.33 (+  7.46) 24.13 (+2.80) 26.13 (+4.80) 26.13 (+4.80)
FARE 56.40 64.20 44.70 18.93 (+5.86) 21.47 (+  8.40) 22.00 (+  8.93) 24.27 (+3.60) 26.27 (+5.60) 25.87 (+5.20)
FDA-$L^{0}$ 58.00 65.90 46.40 13.07 (+2.27) 14.40 (+  3.60) 14.40 (+  3.60) 20.53 (+0.93) 20.53 (+0.93) 20.80 (+1.20)
FDA-$L^{0-1}$ 58.10 66.80 46.10 12.80 (+1.33) 13.33 (+  1.86) 13.33 (+  1.86) 20.67 (+0.27) 19.87 (-0.53) 20.13 (-0.27)
FDA-$L^{all}$ 57.90 65.80 46.40 12.27 (+2.14) 13.20 (+  3.07) 13.47 (+  2.67) 20.13 (-0.67) 21.87 (+1.07) 21.73 (+0.93)
Appendix C Full results for untargeted attacks
In this section, we provide full results for all untargeted attacks on all models and tasks. Specifically, for T2IR and I2TR, results on ALBEF is given in Table.19 and Table.20, and results on BLIP is given in Table.21 and Table.22, respectively. Untargeted attacks for visual grounding on ALBEF are given in Table.23.
Table 19: ASR of white-box untargeted attacks against Text-to-Image Retrieval on Flickr30k. The model is ALBEF. After-attack R@k values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the drop of R@1/5 after attacks.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 95.90 78.54 (21.46) 57.39 (42.61) 74.70 (25.30) 55.89 (44.11) 70.93 (29.07) 48.17 (51.83)
$\nicefrac{{2}}{{255}}$
TeCoA 91.20 81.22 (18.78) 60.20 (39.80) 76.02 (23.98) 58.13 (41.87) 67.95 (32.05) 46.01 (53.99)
TeCoA + FDA-$L^{0-1}$ 91.60 80.73 (19.27) 58.72 (41.28) 68.87 (31.13) 50.20 (49.80) 67.75 (32.25) 44.12 (55.88)
FARE 91.10 74.39 (25.61) 51.52 (48.48) 54.35 (45.65) 70.95 (29.05) 49.29 (50.71) 23.79 (76.21)
FARE + FDA-$L^{0-1}$ 90.60 76.73 (23.27) 53.46 (46.54) 52.95 (47.05) 69.31 (30.69) 49.19 (50.81) 26.32 (73.68)
No Defense 95.90 96.15 (3.85) 92.00 (8.00) 88.11 (11.89) 76.73 (23.27) 87.80 (12.20) 72.97 (27.03)
$\nicefrac{{4}}{{255}}$
TeCoA 91.20 98.98 (1.02) 95.33 (4.67) 85.44 (14.56) 73.91 (26.09) 87.36 (12.64) 71.49 (28.51)
TeCoA + FDA-$L^{0-1}$ 91.60 98.68 (1.32) 94.73 (5.27) 85.09 (14.91) 72.41 (27.59) 87.02 (12.98) 69.47 (30.53)
FARE 91.10 98.08 (1.92) 93.93 (6.07) 80.57 (19.43) 63.56 (36.44) 80.57 (19.43) 63.56 (36.44)
FARE + FDA-$L^{0-1}$ 90.60 98.17 (1.83) 93.90 (6.10) 79.67 (20.33) 62.60 (37.40) 80.18 (19.82) 58.33 (41.67)
Table 20: ASR of white-box untargeted attacks against Image-to-Text Retrieval on Flickr30k. The model is ALBEF. After-attack R@k values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the drop of R@1/5 after attacks.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 85.60 84.85 (15.15) 69.91 (30.09) 77.34 (22.66) 64.58 (35.42) 76.02 (23.98) 57.65 (42.35)
$\nicefrac{{2}}{{255}}$
TeCoA 81.44 87.68 (12.32) 74.26 (25.74) 74.70 (25.30) 61.62 (38.38) 74.49 (25.51) 58.16 (41.84)
TeCoA + FDA-$L^{0-1}$ 81.80 88.59 (11.41) 74.02 (25.98) 73.91 (26.09) 59.89 (40.11) 73.26 (26.74) 56.63 (43.37)
FARE 81.48 84.09 (15.91) 69.48 (30.52) 64.18 (35.82) 74.06 (25.94) 61.80 (38.20) 41.13 (58.87)
FARE + FDA-$L^{0-1}$ 80.14 83.61 (16.39) 68.98 (31.02) 63.70 (36.30) 73.27 (26.73) 61.39 (38.61) 61.25 (58.75)
No Defense 85.60 97.08 (2.92) 93.61 (6.39) 89.11 (10.89) 81.30 (18.70) 88.34 (11.66) 77.89 (22.11)
$\nicefrac{{4}}{{255}}$
TeCoA 81.44 99.13 (0.87) 96.51 (3.49) 86.92 (13.08) 79.35 (20.65) 90.81 (  9.19) 81.08 (18.92)
TeCoA + FDA-$L^{0-1}$ 81.80 99.72 (0.76) 97.61 (2.39) 86.96 (13.04) 78.91 (21.09) 85.82 (14.18) 80.33 (19.67)
FARE 81.48 98.27 (1.73) 95.45 (4.55) 84.96 (15.04) 74.06 (25.94) 84.96 (15.04) 74.06 (25.94)
FARE + FDA-$L^{0-1}$ 80.14 98.35 (1.65) 95.27 (4.73) 83.83 (16.17) 73.27 (26.73) 84.93 (15.07) 70.85 (29.15)
Table 21: ASR of white-box untargeted attacks against Text-to-Image Retrieval on Flickr30k. The model is BLIP. After-attack R@k values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the drop of R@1/5 after attacks.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 97.20 82.45(17.55) 61.79 (38.21) 80.94 (19.06) 67.40 (32.60) 72.22 (27.78) 52.96 (47.04)
$\nicefrac{{2}}{{255}}$
TeCoA 81.50 55.33 (44.70) 32.79 (67.21) 47.73 (52.77) 29.22 (70.78) 44.16 (55.84) 26.40 (73.60)
TeCoA + FDA-$L^{0-1}$ 80.40 51.30 (48.70) 28.74 (71.26) 44.25 (55.75) 28.09 (71.91) 40.56 (59.44) 24.08 (75.92)
FARE 79.40 54.19 (45.81) 28.29 (71.71) 60.72 (+39.28) 36.13 (63.87) 58.65 (41.35) 34.93 (65.07)
FARE + FDA-$L^{0-1}$ 79.30 51.41 (48.59) 29.18 (70.82) 56.62 (+43.38) 33.08 (66.92) 54.01 (45.99) 30.59 (69.41)
No Defense 97.20 99.90 (0.10) 99.60 (  0.40) 95.69 (4.31) 90.57 (9.43) 93.18 (6.82) 83.75 (16.25)
$\nicefrac{{4}}{{255}}$
TeCoA 81.50 96.97 (3.03) 93.94 (  6.06) 80.95 (19.05) 68.40 (31.60) 80.98 (19.02) 66.38 (33.62)
TeCoA + FDA-$L^{0-1}$ 80.40 96.64 (3.36) 92.73 (  7.27) 79.07 (20.93) 65.08 (31.92) 77.99 (22.01) 63.81 (36.19)
FARE 79.40 94.23 (5.77) 84.87 (15.13) 85.09 (14.91) 66.27 (33.73) 83.46 (16.54) 63.76 (36.42)
FARE + FDA-$L^{0-1}$ 79.30 94.14 (5.86) 82.86 (17.14) 82.43 (17.57) 63.12 (36.88) 80.04 (19.96) 60.95 (39.05)
Table 22: ASR of white-box untargeted attacks against Image-to-Text Retrieval on Flickr30k. The model is BLIP. After-attack R@k values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the drop of R@1/5 after attacks.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 87.30 89.68 (10.32) 78.74 (21.26) 85.45 (14.55) 74.51 (25.49) 80.19 (19.81) 65.22 (34.78)
$\nicefrac{{2}}{{255}}$
TeCoA 68.00 62.58 (37.42) 41.35 (58.65) 53.87 (46.13) 34.11 (65.89) 50.67 (49.33) 31.41 (68.59)
TeCoA + FDA-$L^{0-1}$ 67.78 58.08 (41.92) 37.06 (62.94) 50.37 (49.63) 30.10 (69.90) 49.50 (50.50) 28.36 (71.64)
FARE 65.64 63.35 (36.65) 44.21 (55.79) 68.64 (31.36) 48.99 (51.01) 66.25 (33.75) 45.97 (54.03)
FARE + FDA-$L^{0-1}$ 66.22 59.73 (40.27) 39.97 (60.03) 64.06 (35.94) 44.39 (55.61) 61.54 (38.46) 41.74 (58.26)
No Defense 84.02 99.90 (0.10) 99.79 (0.21) 96.70 (  3.30) 93.09 (6.91) 94.01 (5.99) 88.34 (11.66)
$\nicefrac{{4}}{{255}}$
TeCoA 68.00 97.42 (2.58) 93.13 (6.87) 82.94 (17.06) 69.33 (30.67) 80.98 (19.02) 66.38 (33.62)
TeCoA + FDA-$L^{0-1}$ 67.78 96.64 (3.36) 91.04 (8.96) 80.22 (19.78) 64.43 (35.57) 77.99 (22.01) 63.81 (36.19)
FARE 65.64 94.21 (5.79) 88.16 (11.84) 87.41 (12.59) 76.07 (23.93) 86.52 (13.48) 74.31 (25.69)
FARE + FDA-$L^{0-1}$ 66.22 94.45 (5.55) 88.40 (11.60) 86.00 (14.00) 72.89 (27.11) 84.24 (15.76) 71.37 (28.63)
Table 23: Attack success rate (ASR) of untargeted PGD/APGD/MAPGD (masked APGD) against for Visual Grounding (VG) on RefCOCO+. After-attack accuracies are presented in parentheses. All results are in percentage (%). ASR indicates an accuracy drop after attacks.
$l_{\infty}$ Defense Clean Performance Test A Split ($\downarrow$) Test B Split ($\downarrow$)
Val_d Test A Test B PGD APGD MAPGD PGD APGD MAPGD
$\nicefrac{{2}}{{255}}$ No Defense 58.50 65.90 46.30 17.40 (48.50) 15.90 (50.00) 15.30 (50.60) 13.20 (33.10) 7.40 (38.90) 7.60 (38.70)
TeCoA 57.20 64.70 45.00 8.20 (56.50) 9.80 (54.90) 9.80 (54.90) 3.00 (42.00) 4.60 (40.40) 4.70 (40.30)
TeCoA + FDA-$L^{all}$ 57.10 64.90 45.30 8.30 (56.60) 9.80 (55.10) 9.70 (55.20) 3.60 (41.70) 4.80 (40.50) 5.00 (40.30)
FARE 56.40 64.20 44.70 9.40 (54.80) 11.80 (52.40) 12.00 (52.20) 2.90 (41.80) 4.70 (40.00) 4.60 (40.10)
FARE + FDA-$L^{all}$ 56.10 63.70 44.70 9.50 (54.50) 10.90 (53.10) 10.80 (53.20) 3.40 (41.00) 4.00 (40.40) 4.10 (40.30)
$\nicefrac{{4}}{{255}}$ No Defense 58.50 65.90 46.30 21.40 (44.50) 18.70 (47.20) 18.20 (47.70) 14.90 (31.40) 8.60 (37.70) 8.80 (37.50)
TeCoA 57.20 64.70 45.00 8.30 (56.40) 12.50 (52.20) 12.20 (52.50) 2.90 (42.10) 5.90 (39.10) 6.30 (38.70)
TeCoA + FDA-$L^{all}$ 57.10 64.90 45.30 7.90 (57.00) 11.30 (53.60) 11.60 (53.30) 3.40 (41.90) 5.90 (39.40) 6.10 (39.20)
FARE 56.40 64.20 44.70 9.40 (54.80) 11.80 (52.40) 13.60 (50.60) 3.10 (41.60) 5.60 (39.10) 5.70 (39.00)
FARE + FDA-$L^{all}$ 56.10 63.70 44.70 9.50 (54.50) 10.90 (53.10) 12.20 (51.80) 2.70 (41.70) 5.10 (39.30) 5.30 (39.10)
Appendix D Full results for ablation studies
In this section, we provide full results for all ablation studies. T2IR and I2TR results are given in Table.24 and Table.25. Zero-shot performance is given in Table.26.
Table 24: ASR of ablation studies against Text-to-Image Retrieval on Flickr30k. The model is ALBEF. Changes over unattacked values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the attack success rate of the adversarial image showing up in the top-1/5 position of the targeted text queries.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 95.90 0.30 (+0.20) 7.50 (+6.50) 14.60 (+14.50) 15.70 (+14.70) 50.10 (+50.00) 81.90 (+80.90)
$\nicefrac{{2}}{{255}}$
FDA - $\mathcal{T}$ 95.10 0.40 (+0.40) 6.40 (+6.20) 20.70 (+20.70) 23.70 (+23.50)   4.80 (+  4.80) 28.30 (+28.10)
FDA - $\mathcal{T}$ & $\mathcal{H}$ 93.80 0.50 (+0.50) 6.90 (+6.60) 16.80 (+20.90) 19.30 (+19.30) 13.60 (+13.60) 21.80 (+21.80)
FDA - $\mathcal{H}$ 95.60 0.10 (+0.10) 7.30 (+6.60) 12.10 (+12.10) 13.40 (+12.70) 43.60 (+43.60) 73.90 (+73.20)
$L^{all},H^{all}$ 95.50 0.10 (+0.10) 7.30 (+6.80) 14.40 (+14.40) 15.90 (+15.40) 46.50 (+46.50) 84.90 (+84.40)
$L^{all},H^{6-11}$ 95.00 0.20 (+0.20) 8.00 (+7.70) 16.70 (+16.70) 19.50 (+19.20) 48.90 (+48.90) 85.60 (+85.30)
$L^{all},H^{0-5}$ 95.40 0.40 (+0.40) 5.90 (+5.20) 12.80 (+12.80) 14.20 (+13.50) 43.50 (+43.50) 77.30 (+76.60)
$L^{0},H^{0-5}$ 95.60 0.10 (+0.10) 7.30 (+6.60) 12.10 (+12.10) 13.40 (+12.70) 43.60 (+43.60) 73.90 (+73.20)
$L^{0-1},H^{0-5}$ 95.40 0.20 (+0.20) 6.90 (+6.10) 12.00 (+12.00) 12.80 (+12.00) 43.30 (+43.30) 73.60 (+72.80)
Full Dict 95.40 0.40 (+0.40) 5.90 (+5.20) 12.80 (+12.80) 14.20 (+13.50) 43.60 (+43.60) 77.50 (+77.00)
Shortlisted Dict 95.10 0.30 (+0.30) 6.60 (+6.10) 12.80 (+12.80) 14.20 (+13.50) 43.50 (+43.50) 77.30 (+76.60)
No Defense 95.90 4.30 (+4.20) 14.10 (+13.10) 16.50 (+16.40) 16.60 (+15.60) 75.00 (+74.90) 87.00 (+86.00)
$\nicefrac{{4}}{{255}}$
FDA - $\mathcal{T}$ 95.10 5.10 (+0.20) 20.00 (+19.80) 24.80 (+24.80) 25.50 (+25.30) 12.10 (+12.10) 40.00 (+39.80)
FDA - $\mathcal{T}$ & $\mathcal{H}$ 93.80 4.70 (+4.70) 16.70 (+16.40) 20.90 (+20.90) 22.20 (+21.90) 19.90 (+19.90) 23.50 (+23.20)
FDA - $\mathcal{H}$ 95.60 2.90 (+2.90) 13.50 (+12.80) 13.90 (+13.90) 14.10 (+13.40) 69.00 (+69.00) 82.40 (+81.70)
$L^{all},H^{all}$ 95.50 3.90 (+3.90) 14.70 (+14.70) 16.30 (+16.30) 16.70 (+16.20) 76.70 (+76.70) 92.10 (+91.60)
$L^{all},H^{6-11}$ 95.00 3.30 (+3.30) 15.70 (+15.40) 19.60 (+19.60) 20.10 (+19.80) 77.70 (+77.70) 91.70 (+91.40)
$L^{all},H^{0-5}$ 95.40 3.00 (+3.00) 13.50 (+12.80) 14.60 (+14.60) 14.80 (+14.10) 68.90 (+68.90) 84.10 (+83.40)
$L^{0},H^{0-5}$ 95.60 2.90 (+2.90) 13.50 (+12.80) 13.90 (+13.90) 14.10 (+13.40) 69.00 (+69.00) 82.40 (+81.70)
$L^{0-1},H^{0-5}$ 95.40 3.00 (+3.00) 12.40 (+11.60) 13.90 (+13.90) 14.00 (+13.20) 68.10 (+68.10) 81.30 (+80.50)
Full Dict 95.40 3.00 (+3.00) 13.50 (+12.80) 14.90 (+14.90) 15.20 (+14.70) 69.60 (+69.60) 83.90 (+83.40)
Shortlisted Dict 95.10 3.70 (+3.70) 11.60 (+11.10) 12.80 (+12.80) 14.20 (+13.50) 43.50 (+43.50) 77.30 (+76.60)
Table 25: ASR of ablation studies against Image-to-Text Retrieval on Flickr30k. The model is ALBEF. Changes over unattacked values are presented in parentheses. All results are in percentage (%). ASR@1/5 indicates the attack success rate of the adversarial image showing up in the top-1/5 position of the targeted text queries.
$l_{\infty}$ Defense Clean $\uparrow$ PGD APGD MAPGD
ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$ ASR@1 $\downarrow$ ASR@5 $\downarrow$
No Defense 85.60 0.30 (+0.30) 1.10 (+1.10) 14.40 (+14.50) 15.40 (+15.40) 53.50 (+53.50) 63.50 (+63.50)
$\nicefrac{{2}}{{255}}$
FDA - $\mathcal{T}$ 85.28 0.30 (+0.30) 1.50 (+1.50) 20.90 (+20.90) 22.30 (+22.30)   5.90 (+  5.90) 10.40 (+10.40)
FDA - $\mathcal{T}$ & $\mathcal{H}$ 93.80 0.10 (+0.10) 1.10 (+1.10) 16.80 (+20.90) 17.80 (+17.80) 13.00 (+13.00) 15.10 (+15.10)
FDA - $\mathcal{H}$ 85.50 0.10 (+0.10) 0.60 (+0.60) 12.30 (+12.30) 12.80 (+12.80) 46.50 (+46.50) 56.20 (+56.20)
$L^{all},H^{all}$ 85.54 0.30 (+0.30) 1.00 (+1.00) 14.40 (+14.40) 15.00 (+15.00) 51.70 (+51.70) 60.90 (+60.90)
$L^{all},H^{6-11}$ 84.96 0.30 (+0.30) 1.10 (+1.10) 17.30 (+17.30) 17.80 (+17.80) 55.60 (+48.90) 72.30 (+72.30)
$L^{all},H^{0-5}$ 85.40 0.20 (+0.20) 1.00 (+1.00) 13.50 (+13.50) 13.70 (+13.70) 48.50 (+48.50) 58.00 (+58.00)
$L^{0},H^{0-5}$ 85.50 0.10 (+0.10) 0.60 (+0.60) 12.30 (+12.30) 12.80 (+12.80) 46.50 (+46.50) 56.20 (+56.20)
$L^{0-1},H^{0-5}$ 85.32 0.10 (+0.10) 0.80 (+0.80) 12.10 (+12.10) 12.50 (+12.50) 46.90 (+46.90) 55.90 (+55.90)
$L^{all},H^{0-5}$ - Full Dict 84.46 0.40 (+0.40) 5.90 (+5.20) 13.50 (+13.70) 13.70 (+13.70) 48.00 (+48.00) 57.40 (+57.40)
$L^{all},H^{0-5}$ - Shortlisted Dict 85.40 0.20 (+0.20) 1.00 (+1.00) 13.50 (+13.50) 13.70 (+13.70) 48.50 (+48.50) 58.00 (+58.00)
No Defense 85.60 4.50 (+4.50) 9.80 (+9.80) 15.70 (+15.70) 15.90 (+15.90) 74.40 (+74.40) 79.00 (+79.00)
$\nicefrac{{4}}{{255}}$
FDA - $\mathcal{T}$ 95.10 6.40 (+6.40) 11.30 (+11.30) 24.50 (+24.50) 24.90 (+24.90) 15.50 (+15.50) 19.10 (+19.10)
FDA - $\mathcal{T}$ & $\mathcal{H}$ 93.80 5.10 (+5.10) 10.00 (+10.00) 20.80 (+20.80) 21.10 (+21.10) 19.80 (+19.80) 21.00 (+21.00)
FDA - $\mathcal{H}$ 85.50 3.30 (+3.30)   6.50 (+  6.50) 13.70 (+13.70) 13.70 (+13.70) 68.80 (+68.80) 72.40 (+72.40)
$L^{all},H^{all}$ 85.54 5.20 (+5.20) 7.90 (+7.90) 15.90 (+15.90) 16.10 (+16.10) 78.90 (+78.90) 82.50 (+82.50)
$L^{all},H^{6-11}$ 84.96 4.10 (+4.10) 8.80 (+8.80) 19.10 (+19.10) 19.30 (+19.30) 82.00 (+82.00) 85.40 (+85.40)
$L^{all},H^{0-5}$ 85.40 4.00 (+4.00) 7.80 (+7.80) 14.10 (+14.10) 14.20 (+14.20) 72.20 (+72.20) 75.20 (+75.20)
$L^{0},H^{0-5}$ 85.50 3.30 (+3.30) 6.50 (+6.50) 13.70 (+13.70) 13.70 (+13.70) 68.80 (+68.80) 72.40 (+72.40)
$L^{0-1},H^{0-5}$ 85.32 3.10 (+3.10) 6.90 (+6.90) 13.40 (+13.40) 13.50 (+13.50) 69.30 (+69.30) 72.30 (+72.30)
$L^{all},H^{0-5}$ - Full Dict 84.46 3.60 (+3.60) 7.40 (+7.40) 14.10 (+14.10) 14.10 (+14.10) 70.80 (+70.80) 74.40 (+74.40)
$L^{all},H^{0-5}$ - Shortlisted Dict 85.40 4.00 (+4.00) 7.80 (+7.80) 14.10 (+14.10) 14.20 (+14.20) 72.20 (+72.20) 75.20 (+75.20)
Table 26: Zero-shot performance by applying FDA as a plug-and-play tool on T2IR, I2TR on ALBEF/BLIP and VG on ALBEF.
Tasks Models Method Zero-shot Performance($\uparrow$) Average
T2IR/I2TR ALBEF w/o FDA 88.50 98.50 99.20 75.88 93.34 88.50 92.01 -
$L^{all}$ 89.10 98.60 99.40 75.56 93.70 96.66 92.17 $\uparrow$   0.16
$L^{0}$ 89.00 98.50 99.30 75.38 93.20 96.72 92.02 $\uparrow$   0.01
$L^{0-1}$ 89.60 98.80 99.40 76.16 93.70 96.80 92.41 $\uparrow$   0.40
BLIP w/o FDA 87.20 98.00 99.10 78.20 94.08 96.88 92.24 -
$L^{all}$ 88.70 98.40 99.30 78.80 94.20 96.88 92.71 $\uparrow$   0.47
$L^{0}$ 87.00 98.00 99.10 78.14 94.10 96.82 92.19 $\downarrow$   0.05
$L^{0-1}$ 87.10 98.00 99.10 78.12 94.16 96.82 92.22 $\downarrow$   0.02
VG ALBEF w/o FDA 54.50 61.77 43.10 53.12 -
$L^{all}$ 54.73 62.17 43.11 53.34 $\uparrow$   0.22
$L^{0}$ 54.10 61.71 42.34 52.72 $\downarrow$   0.40
$L^{0-1}$ 54.14 61.47 42.42 52.68 $\downarrow$   0.44
Appendix E Details for function word dictionary
We provide the function word dictionary we used as follows: “am, is, are, was, were, be, been, being, have, has, had, do, does, did, will, would, shall, should, may, might, must, can, could, ought, dare, need, used, to, a, an, the, and, but, if, or, because, as, until, while, of, at, by, for, with, about, against, between, into, through, during, before, after, above, below, to, from, in, out, on, off, over, under, again, further, then, once, here, there, when, where, why, how, all, any, both, each, few, more, most, other, some, such, no, nor, not, only, own, same, so, than, too, very”.
Appendix F Visualization of attention scores
Finally, we provide an illustration of original attention, FDA with one subtraction, and FDA, as shown in Fig.4.
Figure 4: A heatmap of attention probabilities given the same image and text inputs. Left: Original attention probabilities are relatively ‘noisy’ and have several visible stripes with very low probabilities, implying the existence of some less relevant visual tokens that are activated, with negligible contributions. Mid: Attention probabilities with one FDA subtraction show much less aforementioned ‘stripes’, with much cleaner and more focused attentions. However, some distractions still exist and remain visible. Right: Attention probabilities with two subtractions show the cleanest attention maps and have the most negligible distractions, with only strong activations on the most relevant visual tokens, i.e., with higher probabilities.